OpenStudy (rock_mit182):

• Present your final solutions as decimal approximations carried out to the third decimal place. Due to the nature of these solutions, no check is required.

OpenStudy (rock_mit182):

\[\frac{ \sqrt{15} }{ 5 }\]

OpenStudy (rock_mit182):

@dan815

OpenStudy (rock_mit182):

please help me out

OpenStudy (rock_mit182):

\[x =\frac{ \pm 2\sqrt{3}\sqrt{5} }{ 2*5 }\]

OpenStudy (rock_mit182):

\[x = \frac{ \pm \sqrt{15} }{ 5 }\]

OpenStudy (rock_mit182):

\[\Delta = 0^{2} - 4(5)(-3)\]

OpenStudy (rock_mit182):

\[\Delta = 60\]